OpenStudy (anonymous):

How would I do this problem √112 + √63 ?

OpenStudy (anonymous):

first do the radicals

OpenStudy (anonymous):

I'm extremely confused.. So I would find the sqrt of 112 and 63? Or...?

OpenStudy (zzr0ck3r):

\[\sqrt{112}=\sqrt{16*7}=\sqrt{16}\sqrt{7}=4\sqrt{7}\\\sqrt{63}=\sqrt{7*9}=\sqrt{9}\sqrt{7}=3\sqrt{7}\\\sqrt{112}+\sqrt{63}=4\sqrt{7}+3\sqrt{7}=?\]

OpenStudy (zzr0ck3r):

hint 4apples + 3apples = 7apples apples = sqrt(7)

OpenStudy (anonymous):

Thanks. Haha. My Brain's kinda gotten fried. So I end up with 7√7, right.

OpenStudy (zzr0ck3r):

correct